Solve each of the following systems by using either addition or substitution. If a unique solution does not exist, state whether the system is dependent or inconsistent.
10x + 2y = 7
y=-5x + 3

Answer by SHUgrad05(58)

10x + 2y = 7
y=-5x + 3
Using Substitution:
10x+2(-5x+3)=7
10x-10x+6=7
6=7
6 does not equal 7 so the system is inconsistent because there is not a solution that exists.